OpenStudy (anonymous):

Write the equation in slope intercept form with a slope of -3 and contains (1,1)

OpenStudy (anonymous):

The equation of a line is \[y = mx + b\] What does each variable in that equation represent?

OpenStudy (anonymous):

I don't know I need help

OpenStudy (anonymous):

y and x are a coordinate pair. For every x, there is exactly 1 y. m is the slope of the line. b is the y-intercept of our line, or the point where our line crosses the y-axis. Since we are given the slope and one point on the line, how could you about solving for b?

OpenStudy (anonymous):

you can also use the point-slope formula, just plug-in the given values to the formula...\[(y-y_1)=m(x-x_1)\]this will yield\[(y-1)=-3(x-1)\]... you can continue finishing it to find the simplified equation...